PEGCOR has a​ P/E ratio of 18.1.Earnings per share are $ 2.72​, and the expected EPS years from today are $ 4.85. Calculate the PEG ratio.
Answer #1

growth rate=(4.85/2.72)^(1/5)-1=12.3%

Calculate the PEG ratio

=(P/E)/growth rate

=(18.1)/12.3

=1.47

the above is answer..